Rationalize the denominator and simplify.
12
----
√7+2

Answer by lwsshak3(11628)

Rationalize the denominator and simplify.
12
----
√7+2
**
12/(√7+2)
multiply top and bottom by the conjugate (√7-2) of the denominator
12(√7-2)/(√7+2)(√7-2)
(12√7-24)/(7-4)
(12√7-24)/(3)
4√7-8
..
note:(7-4)=(√7^2-2^2)=(√7+2)(√7-2)=difference of two squares